OpenStudy (drmoss):

find the 50th term of the sequence 5 -2 -9 -16 ...

OpenStudy (evoker):

Looks like you are subtracting the same number each time.

OpenStudy (drmoss):

yeah its -7 so whats the 50th term?

OpenStudy (drmoss):

how do you create a formula for these sequences

OpenStudy (drmoss):

I give medals ;(

OpenStudy (drmoss):

i c u sam

OpenStudy (drmoss):

halp ;-;

sam (.sam.):

It's an arithmetic progression, what do you think?

sam (.sam.):

To find the 50th term or any term in an AP, firstly, you should find the common difference d. That is you'd pick a number from that sequence and subtract the previous number. Common difference d=-2-5=-7

sam (.sam.):

You can verify your common difference by doing the same for the subsequent numbers, Common difference d=-9-(-2)=-9+2=-7 Now using the AP formula \(a_n=a_1+(n-1)d\) where \(a_n\) is the nth term and \(a_1\) is the first term. \[a_n=a_1+(n-1)d \\ \\ a_{50}=5+(50-1)(-7) \\ \\ a_{50}=?\]

sam (.sam.):

Everything good?

OpenStudy (mathmale):

\[a_n=a_1+(n-1)d \\ \\ a_{50}=5+(50-1)(-7) \\ \\ \]

OpenStudy (mathmale):

Please simplify that quantity within parentheses: (50-1). Your answer? Re-write the expression for a_50 using this result. Share it here, please.

OpenStudy (drmoss):

oh i finished this like an hour ago lol.